OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. y - 2 = 4(x + 6) y + 6 = 4(x - 2) y + 2 = 4(x - 6) y - 6 = 4(x + 2)

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

y-6 = 4(x+2) ; the equation is in point slope form, y-y1 = m(x-x1), where (x1, y1) is the point given and m is the slope of the line

OpenStudy (anonymous):

thxs